Given condition: k1=k2=k3=k4=k5=k,F3=P, and nodes 1,2,5 are fixed
Decision
a.Global stiffness matrix
b. Displacement of nodes 3 and 4
c. Reaction force at nodes 1,2,5
d.Internal force of spring 2,4,51.7 한 스프링계가 아래와 같다. k2 ki -F3 Amino k4 K3 X 주어진 조건: k = k = kg = k 4 = ks = k, F = P, 그리고 절점 1, 2, 5는 고정됨; 결정할 사항: a. 전역 강성행렬

Oh, and I would appreciate it if you could tell the answer using computer typing instead of handwriting.

Because I don't know the cursive

If it is difficult, I would appreciate it if you could respond in a convenient way.
Answer #1

Given : k2 M F3 K3 ko 3 3. А UT 4 КА 2 2 Ki = K2=kz=Khak5 =k F = P. solution : for element ☺ KI -ki A ki U3 Fз -ki for elemenfor element 4 KE _k5 -k5 ko us Global stiffness matix : ki OK k4 0 -K4 o o -ki kitkatk3 -k2-k3 0 -ky -k2-k3 Katkzt kytk5 -k50 UD A K (FO U12.00 6 -k 6 0 K F2=0 uz F3 =P 3k -2k o -k - ak 4k -k 14 6 FHEO -F5=0 ak k u5=0 6 0 o the rows and columns of Bdi Reaction force : {R} - [K] {u} _{f} ~KU3 -0 => -Klak R,= ald O-k R2 = kuz - ku4 = ,4K R2= - P 4 R5 -* (*) -0 = -kly-ku5= -